Bill has been accepted into a university and is looking into his housing options.   He is considering purchasing a mobile home to live in for the 4 years he will be going to school.   The initial purchase price for the mobile home is $40,000.   Luckily, Bill knows two friends from high school who are willing to be his roommates and pay $400 per month, each.   Bill figures that his lot rent will be $270 per month and taxes, utilities and insurance will be another $300 per month as well.   Bill’s roommates will not pay utilities. Also, by buying the home, Bill will save $7,200 per year in rent, adding to his net returns.   After the four years, Bill hopes to sell the home for $20,000.   Assume straight-line depreciation over 7 years and a marginal tax rate of 20%. Bill requires a pre-tax rate of return of 10%.

What is the after-tax terminal value?

$20,000                                      

$19,428.57

$35,000                                       

$19,000
Answer #1

Annual Depreciation=Cost of Asset/Useful Life=40000/7=$5714.29

Accumulated Depreciation at the end of four years=4*5714.29=$22,857.14

Book value at the end of four years=40000-22857.14=$17,142.86

Sales Price at the end of four years =$20,000

Gain on sale of mobile home at end of year4=20000-17142.86=$2,857.14

Tax Rate=20%

Tax on gain on sale=20%*2857.14=$571.43

After tax terminal value =20000-571.43=$19,428.57

ANSWER:

$19,428.57